OpenStudy (anonymous):

find the lateral area of a cylinder and the total area of a cylinder if r=6cm and h=15cm

OpenStudy (anonymous):

Lateral Surface area = 2πrh = 2*3.14*6 cm * 15 cm = 565.2 cm^2 Total surface Area = 2πrh + 2πr^2 = 2πr[h + r] = 2*3.14*6cm[15cm + 6cm] = 2*3.14*6cm*21cm = 791.28 cm^2